# Common origin of black holes in high mass X–ray binaries and in gravitational–wave sources

### v3

@@ 5 Cyg X-1 spin from the soft component
We confront two models of Cyg X-1 emission. We use data from the softest state ever seen (observed by Suzaku in 2013; magenta star in Fig. 2), when the soft component is at its most dominant, and so uncertainties from modelling the high energy coronal tail are smallest (observation B4 from Kawano et al. 2017, see this paper for details of the data). We fit these data using the xspec spectral fitting package and all the models below include absorption along the line of sight using tbabs with column density as a free parameter (see Appendix A for the details). We show results from a series of model fits in Figure 3, where the original (absorbed) X-ray Suzaku data is grey, with the reconstructed (corrected for absorption) spectrum shown with orange points. We assume here the standard mass and distance, and fix the disc inclination to the binary inclination of .
@@
We first fit the standard disc model to these data. The results are shown in the left panel of Figure 3. This includes all special and general relativistic ray tracing effects, as well as allowing a correction to the blackbody temperatures to account for the incomplete thermalisation expected from full models of the vertical disc structure (colour temperature correction, fixed to ). We allow a fraction of these disc spectra to be Compton upscattered to form the power law tail, modelled using the standard formalism, which assumes the electron energy is typically higher than keV. We recover the usual, very high black hole spin () seen in all standard-disc fits to the soft component (Zhao et al. 2021a; Zdziarski et al. 2024a, b).
@@
The middle panel of Figure 3 shows what happens when we additionally allow a warm skin on the top of the disc. We model this using another Compton upscattering model, which allows the electron temperature to be a free parameter, as well as the scattered fraction (set by the optical depth of the skin). This is formally a much better fit to the data, even accounting for the additional free parameters: optical depth and temperature of the warm disc skin. A simple F-test gives a probability of the improvement being random as . This is well below the generally accepted limit of for a result to be significant. More importantly, the best fit reduces BH spin to . The BH spin here is extremely poorly constrained as the soft component shape in this model is set more by the electron temperature of the skin than by the temperature of the disc at the ISCO. We illustrate this in the right panel of Figure 3, where we show a fit with almost the same but for , as predicted by standard stellar evolution.
@@
This demonstrates that formal solutions for BH spin from spectra with their reported extremely small error bars or very definitive values (e.g., at level (Zhao et al. 2021b)) are model-dependent and are not the only solution (see also Zdziarski et al. 2024a, b who arrive at the same conclusion).
5.1 Can a disc form in Cyg X-1 for low BH spin?
@@
Modelling of the disc continuum spectra assumes that an accretion disc can form around the BH for any value of the BH spin and/or any orientation of the accretion disc with respect to the spin of the BH. In a binary system, a disc will form only if the circularisation radius of the transferred matter is larger than the radius of the accreting body. In other words, a disc can form if the angular momentum of the accreting matter is large enough to allow it to circle around the accretor. For wind (Bondi-Hoyle; Bondi 1952) accretion the condition is (Frank et al. 2002)
@@
where is the mass of the donor star, the orbital separation, the wind velocity and the escape velocity from the donor. For the parameters of Cyg X-1, when , cm but cm, so that e.g. for no disc can form. Sen et al. (2021) analysed the problem in great detail and showed that for the observed binary and stellar parameters of Cyg X-1 (Miller-Jones et al. 2021), a maximally spinning BH may be required to form an accretion disc in the prograde orientation when the BH accretes from a wind.
@@
However, we note that the O star companion in Cyg X-1 is nearly filling its Roche lobe (Roche-lobe filling factor is greater than 0.9, Miller-Jones et al. 2021), where the formation of a focused accretion stream is expected (Blondin et al. 1991; Hadrava & Čechura 2012; El Mellah et al. 2019). In Cyg X-1, the existence of a focused accretion stream has also been observationally verified (Miller et al. 2005; Poutanen et al. 2008; Hanke et al. 2009). A high Roche lobe filling factor has been found for LMC X-1 (Orosz et al. 2009) and the O donor in M33 X-7 effectively fills its Roche lobe (although not in a standard way, see Ramachandran et al. 2022).
@@
Hence, the criterion of the formation of an accretion disc in Cyg X-1 does not rule out any assumptions on the spin of the BH. The high Roche lobe filling factor of the O star enables the formation of a focused wind with sufficient angular momentum to make a prograde accretion disc. In such case the circularisation radius for the bulk fo the accretion flow is of the order of the solar radius (see, e.g., Frank et al. 2002), orders of magnitude larger than the ISCO radius. The above arguments, together with our modelling of the disc continuum spectra imply that the observations of Cyg X-1 are consistent with the statement that its BH can also be slowly spinning.
@@ 6 M33 X-7 BH spin from the soft component
Assuming the older BH mass, Liu et al. (2008) used the continuum fitting method to get a BH spin . We have used the standard continuum method to determine the M33 X-7 spin value for the new BH mass and obtained (see Appendix B), as expected. With this new spin value, the M33 X-7 BH is no longer in the very-rapidly rotating class, to which Cyg X-1 with and LMC X-1 with belong. Nevertheless, since we have concluded above that the spin determination in Cyg X-1 is strongly model-dependent and Zdziarski et al. (2023, 2024b) arrived to same conclusion regarding LMC X-1, we thought it useful to check if the same applies to M33 X-7. Similarly to these sources, when a warm-skin Comptonizing model is used, the black hole spin is completely unconstrained, with acceptable values ranging from -1 to at confidence level (see Appendix B).
Below, we will discuss the possible reason for this common property of the only three HMBXs that are potential progenitors of BHBH binary systems.
@@ 7 Discussion
We have shown that the X-ray spectra of Cyg X-1 and M88 X-7 can be fitted with both standard disk and warm Comptonised disc models (see also Zdziarski et al. 2024a, b for the same conclusion about Cyg X-1 and LMC X-1). We have also noticed that even in its softest state, the spectrum of Cyg X-1 is not soft enough to justify a standard-disc model fit, so there are very good reasons to consider alternatives.
@@
In order to avoid ambiguities and misunderstandings, one has to define precisely what is understood by fitting spectra with a “standard disc” model. This model is describing an axially symmetric, stationary, geometrically thin Keplerian accretion flow. It is assumed that the angular momentum transport is provided by a “viscous” (in astrophysical discs, the viscosity must be anomalous) torque and that the local dissipation of the same torque is the source of the energy radiated away from the disc surface. Then, from mass and angular momentum conservation equations for the vertically averaged disc parameters, assuming thermal equilibrium, one obtains
@@
where is the effective temperature and contains information about the boundary conditions. This is the equation that provides the flux, or effective temperature profile for a Keplerian stationary disc. It does not say that the disc emits radiation like a black body. It contains no information about the vertical energy transport in the disc and is independent of the angular momentum and dissipation mechanism it only assumes the existence of such mechanism (it is independent of the value of the “viscosity parameter , it does not even assume the existence of such a parameter). Eq. (3) is the basis of the “standard” disc X-ray spectrum model. There, one assumes that the colour temperature is close to the effective temperature. i.e., that the disc, locally radiates approximately like a black body. Since the accretor is a BH, one adds to the model general-relativistic effects, including radiation back-falling onto the disc, and appropriate inner boundary conditions (usually zero torque at ISCO). Since the observed spectra are not strictly a sum of black-bodies, a disc annulus emission is represented by a colour-corrected (or diluted) black-body
@@
where is the intensity, is the Planck spectrum, the Planck constant, the radiation frequency and is a colour temperature correction factor representing the extent to which electron scattering through the disc vertical structure is expected to affect the emergent spectrum. This can be calculated by solving the disc vertical radiative transfer equations using a code such as TLUSTY (see Davis & El-Abd 2019; Davis et al. 2005 and references therein). However, contrary to to stellar atmospheres, accretion discs are not in radiative equilibrium, i.e. the divergence of the radiative flux is not equal to zero, but to the (viscous) heating rate per unit volume. Hence the calculations requires the viscous–dissipation vertical structure as input. In the -model of Shakura & Sunyaev (1973), one assumes that this dissipation profile is identical to that of the total pressure, but this is radiation pressure unstable at the observed disc luminosities. Instead, for the disc calculations, the dissipation is assumed to follow the density profile in TLUSTY Done & Davis (2008). Both these prescriptions have maximal dissipation at the mid-plane, whereas MRI simulations often show maximum dissipation closer to the disc surface (see, e.g., Blaes 2014). Some versions of TLUSTY (Hubeny & Hubeny 1998) use a step-wise power-law dependence of viscosity that allows a chosen fraction of the dissipation to occur in the surface layers. Such dissipation profile is needed to reproduce disc spectra of cataclysmic variables Hubeny & Long (2021) and when applied to BH X-ray emitting disc, such dissipation stratification produces disc coronas (Davis & El-Abd 2019) as suggested by Svensson & Zdziarski (1994).
@@
What is, or rather what are the real, physical dissipation profiles is still unknown since numerical simulations cannot yet produce realistic vertical structures of geometrically thin accretion discs. That is why, disc models with warm surface layers are as physically legitimate as the presumed “standard” model, especially when there is strong evidence as in Cyg X-1 that the structure is not a standard disc.
@@
One should also point out that the -disc model cannot explain the brightness–hardness ratio tortoise-head hysteresis observed in transient X-ray sources. The physical reason for this variability is still unknown. Observations (e.g. Done et al. 2007; Basak & Zdziarski 2016; Basak et al. 2017) and models (e.g., Dubus et al. 2001) suggests that accretion discs in BH X-ray binaries are truncated, but the physical mechanism of this excision is still a mystery.
@@
It does not mean that the -disc model on which the “standard” spectral model is based, is to be rejected. This model has found many successful applications. The soft state spectra in transients show peak disk temperature and total luminosity changing together with , indicating a constant inner radius for the disc and constant Done et al. (2007). Even the radial dependence of the disc temperature shows the predicted behaviour in eclipsing cataclysmic variables in their disk (outbursting) states (Horne & Cook 1985), as well as showing the very different radial profile expected from non-equilibrium discs in quiescent dwarf-novae; Lasota 2001.
@@
None of the known HMBH show a convincing soft state relation which characterises a standard disc (see Sect. 4). Cyg X-1 is never in a soft state, while in LMC X-1, Gierliński & Done (2004a) discovered that the disc luminosity follows instead a track. Zdziarski et al. (2024a) found that such a anti-correlation between luminosity and temperature can be naturally explained as an effect of the warm Comptonization layer on top of a standard disc. The luminosity variations in M88 X-7 are not large enough to determine whether a relation is present 222Liu et al. (2008) present a variability in counts (not in counts per second) which reflects mostly the variability of the exposure time..
@@
The three BHHMBXs, Cyg X-1, LMC X-1 and M33 X-7, whose spectra are well fitted by a model of a disc covered with warm comptonising layer have one feature in common: the accrete not only from focused stream but also capture some matter from the strong wind of the BH companion. It is natural to speculate that this is the source of the warm layer, making it a ubiquitous feature of the HMBH spectra.
@@ 8 Conclusion
The difference between HMXB and LVK BH masses has already been noticed after the first gravitational wave detection and attributed to metallicity effect (Abbott et al. 2016) as anticipated by Belczynski et al. (2010b). Taking into account that interferometric detectors of gravitational waves sample a volume that is a few hundred thousand times larger than the X-ray telescopes (for HMXB BH mass determination), the difference in BH masses between LVK and HMXBs is not surprising as it implies a huge difference in the metallicity range sampled (see Fig. 1).
@@
While assessing the problem of BH mass difference between HMXB and gravitational–wave sources is rather straightforward, addressing the apparent conflicting spin values is more complicated. At face value, the spin values attributed to the BH in the HMXBs that are massive enough to form LVK sources (Cyg X- 1, LMC X-1, M33 X-7) seem to require two distinct populations: slowly spinning (LVK) and rapidly spinning (HMXBs) BHs.
@@
Since the majority of LVK BHs are slowly spinning (), we argue, under the assumption that they are stellar-origin BHs (i.e., are not primordial BHs: Hawking 1971, but see Mroz et al. 2024), that it must indicate that massive stars are subject to efficient angular momentum transport (Spruit 2002b; Fuller et al. 2019; Eggenberger et al. 2022; Petitdemange et al. 2024). This leads to good agreement of predicted BH spins in BH-BH mergers with LVK data (Belczynski et al. 2020b; Bavera et al. 2020). These slowly spinning BHs can form in isolated binary evolution and various dense (open, globular, nuclear) cluster environments. A small fraction of LVK BHs may be moderately or even rapidly spinning (see the long tail of BH spins with : left panel of Fig.15 in The LIGO Scientific Collaboration et al. 2023). These moderately/rapidly spinning BHs can also be explained in the framework of efficient angular momentum transport. Although the star rotation slows down during its evolution towards the formation of a BH, the tidal spin-up of a WR star (BH progenitor) in close BH+WR binary systems (BH-BH progenitors) can easily result in a moderately or rapidly spinning BH (Olejak & Belczynski 2021). In dense clusters, dynamical interactions may lead to the formation of BH-BH mergers with components that are second or even third-generation BHs (formed by earlier BH-BH mergers) that naturally have significant spins (Rodriguez et al. 2019).
@@
Under the natural assumption that BHs in HMXBs are formed directly from massive stars and within the paradigm of efficient angular momentum transport, these BHs also must have low spins (; e.g., Belczynski et al. 2020b). We have demonstrated that this conclusion is fully consistent with BH spin estimate based on fitting the disc spectra in HMXBs with models in which the disc is covered by a warm Comptonized layer. In particular, we show that with this model, the X-ray spectra of the softest, but still pretty hard, state of Cyg X-1 is consistent with a slowly spinning BH as confirmed by Zdziarski et al. (2024b) for another set of data. The same is true of X-7 in M33 whose BH spin is completely unconstrained when its soft X-ray spectra are fitted with the warm–skin model. The same warm–skin solutions also give a low spin for LMC X-1 (Zdziarski et al. 2024a. We have shown that the warm–skin solution is as physically plausible as the so–called standard solution, especially in the case of systems that are not in a generic soft X-ray spectral state. We speculate that this warm skin is ubiquitous in the HMXB-BH due to the presence of additional accretion from lower angular momentum material in the stellar wind as well as Roche Lobe Overflow.

### v4

@@ 5 Cyg X-1 spin from the soft component
We confronted two models of Cyg X-1 emission. We used data from the softest state ever seen (observed by Suzaku in 2013; magenta star in Fig. 2), when the soft component is at its most dominant, and, thus, uncertainties from modelling the high energy coronal tail are smallest (observation B4 from Kawano et al. 2017, see this paper for details of the data). We fit these data using the xspec spectral fitting package and all the models below include absorption along the line of sight using tbabs with column density as a free parameter (see Appendix A for the details). We show results from a series of model fits in Fig. 3, where the original (absorbed) X-ray Suzaku data is grey, with the reconstructed (corrected for absorption) spectrum shown with orange points. We assume here the standard mass and distance, and we fix the disc inclination to the binary inclination of .
@@
We first fit the standard disc model to these data. The results are shown in the left panel of Fig. 3. This includes all special and general relativistic ray tracing effects, as well as allowing a correction to the blackbody temperatures to account for the incomplete thermalisation expected from full models of the vertical disc structure (colour temperature correction, fixed to ). We allowed for a fraction of these disc spectra to be Compton upscattered to form the power law tail, modelled using the standard formalism, which assumes the electron energy is typically higher than keV. We recovered the usual, very high black hole spin () seen in all standard-disc fits to the soft component (Zhao et al. 2021a; Zdziarski et al. 2024a, b).
@@
The middle panel of Fig. 3 shows what happens when we additionally allow a warm skin on the top of the disc. We model this using another Compton upscattering model, which allows the electron temperature to remain as a free parameter, as well as the scattered fraction (set by the optical depth of the skin). This is formally a much better fit to the data, even accounting for the two additional free parameters: optical depth and temperature of the warm disc skin. A simple F-test gives a probability of the improvement being random as . This is well below the generally accepted limit of for a result to be significant. More importantly, the best fit reduces BH spin to . The BH spin here is extremely poorly constrained as the soft component shape in this model is set more by the electron temperature of the skin than by the temperature of the disc at the ISCO. We illustrate this in the right panel of Fig. 3, where we show a fit with almost the same but for , as predicted by standard stellar evolution.
@@
This demonstrates that formal solutions for BH spin from spectra with their reported extremely small error bars or very definitive values (e.g. at level, Zhao et al. 2021b) are model-dependent and are not the only solution (see also Zdziarski et al. 2024a, b who arrive at the same conclusion).
5.1 Considering whether a disc can form in Cyg X-1 for a low BH spin
@@
The modelling of the disc continuum spectra assumes that an accretion disc can form around the BH for any value of the BH spin and/or any orientation of the accretion disc with respect to the spin of the BH. In a binary system, a disc will form only if the circularisation radius of the transferred matter is larger than the radius of the accreting body. In other words, a disc can form if the angular momentum of the accreting matter is large enough to allow it to circle around the accretor. For wind accretion (Bondi-Hoyle; Bondi 1952), the condition is (Frank et al. 2002):
@@
where is the mass of the donor star, the orbital separation, the wind velocity, and the escape velocity from the donor. For the parameters of Cyg X-1, when , cm but cm, so that, for instance, for no disc can form. Sen et al. (2021) analysed the problem in great detail and showed that for the observed binary and stellar parameters of Cyg X-1 (Miller-Jones et al. 2021), a maximally spinning BH may be required to form an accretion disc in the prograde orientation when the BH accretes from a wind.
@@
However, we note that the O star companion in Cyg X-1 is nearly filling its Roche lobe (Roche-lobe filling factor is greater than 0.9, Miller-Jones et al. 2021), where the formation of a focused accretion stream is expected (Blondin et al. 1991; Hadrava & Čechura 2012; El Mellah et al. 2019). In Cyg X-1, the existence of a focused accretion stream has also been observationally verified (Miller et al. 2005; Poutanen et al. 2008; Hanke et al. 2009). A high Roche lobe filling factor has been found for LMC X-1 (Orosz et al. 2009) and the O donor in M33 X-7 effectively fills its Roche lobe (although not in a standard way; see Ramachandran et al. 2022).
@@
Hence, the criterion of the formation of an accretion disc in Cyg X-1 does not rule out any assumptions on the spin of the BH. The high Roche lobe filling factor of the O star enables the formation of a focused wind with sufficient angular momentum to make a prograde accretion disc. In such cases, the circularisation radius for the bulk of the accretion flow is of the order of the solar radius (see e.g. Frank et al. 2002), orders of magnitude larger than the ISCO radius. The above arguments, together with our modelling of the disc continuum spectra, imply that the observations of Cyg X-1 are consistent with the notion that its BH could also be slowly spinning.
@@ 6 M33 X-7 BH spin from the soft component
Assuming the older BH mass, Liu et al. (2008) used the continuum fitting method to get a BH spin . We have used the standard continuum method to determine the M33 X-7 spin value for the new BH mass and obtained (see Appendix B), as expected. With this new spin value, the M33 X-7 BH is no longer in the very-rapidly rotating class, to which Cyg X-1 with and LMC X-1 with belong. Nevertheless, since we have concluded that the spin determination in Cyg X-1 is strongly model-dependent, supported by the findings of Zdziarski et al. (2023, 2024b), who arrived to same conclusion regarding LMC X-1, we thought it useful to check if the same applies to M33 X-7. Similarly to these sources, when a warm-skin Comptonising model is used, the black hole spin is completely unconstrained, with acceptable values ranging from -1 to at confidence level (see Appendix B). Below, we discuss the possible reason for this common property of the only three HMBXs that are potential progenitors of BHBH binary systems.
@@ 7 Discussion
We have shown that the X-ray spectra of Cyg X-1 and M33 X-7 can be fitted with both standard disc and warm Comptonised disc models (see also Zdziarski et al. 2024a, b for the same conclusion about Cyg X-1 and LMC X-1). We have also noticed that even in its softest state, the spectrum of Cyg X-1 is not soft enough to justify a standard-disc model fit, so there are very good reasons to consider alternatives.
@@
In order to avoid ambiguities and misunderstandings, one has to define precisely what is understood by fitting spectra with a ‘standard disc’ model. This model describes an axially symmetric, stationary, geometrically thin Keplerian accretion flow. It is assumed that the angular momentum transport is provided by a ‘viscous’ (noting that in astrophysical discs, the viscosity must be ’anomalous’) torque and that the local dissipation of the same torque is the source of the energy radiated away from the disc surface. Then, from the mass and angular momentum conservation equations for the vertically averaged disc parameters, assuming thermal equilibrium, we obtain
@@
where is the effective temperature and contains information about the boundary conditions. This is the equation that provides the flux – or effective temperature profile, – for a Keplerian stationary disc. It does not say that the disc emits radiation like a blackbody. It contains no information about the vertical energy transport in the disc and is independent of the angular momentum and dissipation mechanism; in fact, it only assumes the existence of such mechanism (independently of the value of the viscosity parameter, , and does not even assume the existence of such a parameter). Equation (3) is the basis of the ‘standard’ disc X-ray spectrum model. There, it is assumed that the colour temperature is close to the effective temperature; namely, the disc, locally radiates approximately as a blackbody would. Since the accretor is a BH, one adds to the model general-relativistic effects, including radiation back-falling onto the disc, and appropriate inner boundary conditions (usually zero torque at ISCO). Since the observed spectra are not strictly a sum of blackbodies, the disc annulus emission is represented by a colour-corrected (or diluted) blackbody:
@@
where is the intensity, is the Planck spectrum, the Planck constant, the radiation frequency, and is a colour temperature correction factor representing the extent to which electron scattering through the disc vertical structure is expected to affect the emergent spectrum. This can be calculated by solving the disc vertical radiative transfer equations using a code such as TLUSTY (see Davis & El-Abd 2019; Davis et al. 2005 and references therein). However, contrary to stellar atmospheres, accretion discs are not in radiative equilibrium, namely, the divergence of the radiative flux is not equal to zero, but to the (viscous) heating rate per unit volume. Hence, the calculations requires the viscous-dissipation vertical structure as input. In a simplistic application of the -model of Shakura & Sunyaev (1973), it is assumed that this dissipation profile is identical to that of the total pressure. Instead, for the disc emission calculations, the dissipation is assumed to follow the density profile in TLUSTY (Done & Davis 2008). Both these prescriptions have maximal dissipation at the mid-plane, whereas MRI simulations often show maximum dissipation closer to the disc surface (see e.g. Blaes 2014). Some versions of TLUSTY (Hubeny & Hubeny 1998) use a step-wise power-law dependence of viscosity that allows for a chosen fraction of the dissipation to occur in the surface layers. Such a dissipation profile is needed to reproduce disc spectra of cataclysmic variables Hubeny & Long (2021) and when applied to BH X-ray emitting disc, such dissipation stratification produces disc coronas (Davis & El-Abd 2019), as suggested by Svensson & Zdziarski (1994).
@@
In this regard, the real, physical dissipation profiles are still unknown since numerical simulations cannot yet produce realistic vertical structures of geometrically thin accretion discs. That is why disc models with warm surface layers are as physically legitimate as the presumed ‘standard’ model, especially when there is strong evidence as in Cyg X-1 that the structure is not a standard disc.
@@
It is also useful to point out that the -disc model cannot explain the brightness-hardness ratio tortoise-head hysteresis observed in transient X-ray sources. The physical reason for this variability is still unknown. Observations (e.g. Done et al. 2007; Basak & Zdziarski 2016; Basak et al. 2017) and models (e.g. Dubus et al. 2001) suggests that accretion discs in BH X-ray binaries are truncated, but the physical mechanism of this excision is still a mystery.
@@
This does not mean that the -disc model on which the ‘standard’ spectral model is based is to be rejected. This model has found many successful applications. The soft state spectra in transients show peak disk temperature and total luminosity changing together with , indicating a constant inner radius for the disc and constant (Done et al. 2007). The radial dependence of the disc temperature shows the predicted behaviour in eclipsing cataclysmic variables in their disk (outbursting) states (Horne & Cook 1985), as well as showing the very different radial profile expected from non-equilibrium discs in quiescent dwarf-novae; (Lasota 2001).
@@
None of the known HMBHs exhibit a convincing soft state relation that characterises a standard disc (see Sect. 4). Cyg X-1 is never in a soft state, while in LMC X-1 Gierliński & Done (2004a) discovered that the disc luminosity follows instead a track. Zdziarski et al. (2024a) found that such a anti-correlation between luminosity and temperature can be naturally explained as an effect of the warm Comptonisation layer on top of a standard disc. The luminosity variations in M33 X-7 are not large enough to determine whether a relation is present 333Liu et al. (2008) present a variability in counts (not in counts per second) which reflects mostly the variability of the exposure time..
@@
The three BHHMBXs, Cyg X-1, LMC X-1, and M33 X-7, whose spectra are well fitted by a model of a disc covered with warm comptonising layer have one feature in common: they accrete not only from a focused stream, but also capture some matter from the strong wind of the BH companion. It is natural to speculate that this is the source of the warm layer, making it a ubiquitous feature of the HMBH spectra.
@@ 8 Conclusion
The difference between HMXB and LVK BH masses has already been noticed after the first gravitational wave detection and attributed to metallicity effect (Abbott et al. 2016), as anticipated by Belczynski et al. (2010b). Taking into account that interferometric detectors of gravitational waves sample a volume that is a few hundred thousand times larger than the X-ray telescopes (for HMXB BH mass determination), the difference in BH masses between LVK and HMXBs is not surprising as it implies a huge difference in the metallicity range sampled (see Fig. 1).
@@
Although assessing the problem of BH mass differences between HMXB and gravitational–wave sources is rather straightforward, the task of addressing the apparent conflicting spin values is more complicated. At face value, the spin values attributed to the BH in the HMXBs that are massive enough to form LVK sources (Cyg X- 1, LMC X-1, M33 X-7) seem to require two distinct populations: slowly spinning (LVK) and rapidly spinning (HMXBs) BHs.
@@
Since the majority of LVK BHs are slowly spinning (), we argue, under the assumption that they are stellar-origin BHs (i.e. they are not primordial BHs: Hawking 1971, but see Mróz et al. 2024), that this must indicate that massive stars are subject to efficient angular momentum transport (Spruit 2002b; Fuller et al. 2019; Eggenberger et al. 2022; Petitdemange et al. 2024). This leads to good agreement of predicted BH spins in BH-BH mergers with LVK data (Belczynski et al. 2020b; Bavera et al. 2020). These slowly spinning BHs can form in isolated binary evolution and various dense (open, globular, nuclear) cluster environments. A small fraction of LVK BHs may be moderately or even rapidly spinning (see the long tail of BH spins with : left panel of Fig.15 in The LIGO Scientific Collaboration et al. 2023). These moderately and rapidly spinning BHs can also be explained in the framework of efficient angular momentum transport. Although the star rotation slows down during its evolution towards the formation of a BH, the tidal spin-up of a WR star (BH progenitor) in close BH+WR binary systems (BH-BH progenitors) can easily result in a moderately or rapidly spinning BH (Olejak & Belczynski 2021). In dense clusters, dynamical interactions may lead to the formation of BH-BH mergers with components that are second- or even third-generation BHs (formed by earlier BH-BH mergers) that naturally exhibit significant spins (Rodriguez et al. 2019).
@@
Under the natural assumption that BHs in HMXBs are formed directly from massive stars and within the paradigm of efficient angular momentum transport, these BHs also must have low spins (; e.g. Belczynski et al. 2020b). We have demonstrated that this conclusion is fully consistent with BH spin estimate based on a fitting of the disc spectra in HMXBs with models where the disc is covered by a warm Comptonised layer. In particular, we show that with this model, the X-ray spectra of the softest (but still pretty hard) state of Cyg X-1 is consistent with a slowly spinning BH, as confirmed by Zdziarski et al. (2024b) for another set of data. The same is true of X-7 in M33, whose BH spin is completely unconstrained when its soft X-ray spectra are fitted with the warm-skin model. The same warm-skin solutions also give a low spin for LMC X-1 (Zdziarski et al. 2024a. We have shown that the warm-skin solution is as physically plausible as the so–called standard solution, especially in the case of systems that are not in a generic soft X-ray spectral state. We speculate that this warm skin is ubiquitous in the HMXB-BH due to the presence of additional accretion from lower angular momentum material in the stellar wind as well as Roche Lobe overflow.